OpenStudy (anonymous):

What is the standard form of the equation of the conic given by 2x^2 + 2y^2 + 4x - 12y - 22 = 0 ?

OpenStudy (anonymous):

2(x^2+2x+1)+2(y^2-6y+3)=30 2(x+1)^2+2(y-sqrt(3))^2=30 (X+1)^2+(y-sqrt(3))^2=15

OpenStudy (anonymous):

Then how do I convert it to standard form?

OpenStudy (anonymous):

Sorry, I don't really understand what you mean by "standard form"... If you can give me an example, then I might try again :D

OpenStudy (anonymous):

like make it equal to 1

OpenStudy (anonymous):

That's simple! You just need to divide both sides of the equation by 15!

OpenStudy (anonymous):

ok thank you!